### v1

@@ Abstract
3D direct numerical simulation of electrolysis is applied to investigate the growth and detachment of bubbles at electrodes. The moving gas-liquid interface is modeled employing the VOF-based method. To ensure the accuracy of the simulations, a mesh-independence study has been performed. The simulations include the growth phase of the bubbles followed by their detachment from the electrode surface and the results are validated with analytical models and experimental data. The bubble growth is diffusion controlled leading to the scaling $R=2\beta t^{1/2}$, but the growth exponent is overpredicted by our simulation during initial stage. Furthermore, it is proved that the nucleation sites of the bubble strongly influence gas transport by measuring the relevant Sherwood number. Finally, we investigate the effects of contact angle and nucleation sites on bubble detachment behavior, and compare the detachment radius with Fritz’s formula, the results show a good agreement, confirming that buoyancy is the dominant driving force. As the nucleation sites increase, the induced bubble coalescence accelerates the bubble detachment. Taken together, these findings give us valuable insights into improving gas bubble removal and enhancing overall electrolysis efficiency.
@@ 1 Introduction
The production of green hydrogen through water electrolysis is expected to be an important technology in achieving global ”Net-zero emission” targets (Turner, 2004; Holladay et al., 2009; Dawood et al., 2020). However, it is slow and inefficient in many situations. The attached bubbles in such electrochemical devices reduce the efficiency of electrolyzer systems by blocking the active electrode sites or by increasing the ohmic resistance (Swiegers et al., 2021). Maintaining a bubble-free electrode surface is therefore crucial for high efficient $\mathrm{H_{2}}$ production. Consequently, a detailed understanding of bubble evolution dynamics is highly desirable for developing strategies to enhance water electrolysis efficiency. Analytical solutions for this problem are not generally available, and only a few exact solutions can be derived using extremely simplified assumptions (Epstein and Plesset, 1950; Scriven, 1959). However, they can still serve as valuable references for experimental and numerical studies (Glas and Westwater, 1964; Brandon and Kelsall, 1985; Dapkus and Sides, 1986; van der Linde et al., 2017; Taqieddin et al., 2017; Soto et al., 2018; Angulo et al., 2020; Zhang and Lohse, 2023).
@@
Several experiments have been done, and they confirmed that bubble growth is mainly driven by mass transport due to the gradient of the dissolved hydrogen concentration. Glas and Westwater (1964) performed a simple experiment for hydrogen bubbles generated on a flat electrode. The experimental findings demonstrate that the asymptotic growth of electrochemically generated bubbles follows the same functional relationship given by the analytical solution for a suspended bubble growing in a supersaturated liquid, see Scriven (1959). Recent studies have already investigated this problem from different scales. Macroscopic processes are mainly convection phenomena in the electrolyte solution, which is observed at the electrolytic cell scale. Li et al. (2018) visualized experimentally on two-phase flow at the anode side of a proton exchange membrane electrolyzer. They understood that the inlet velocity does not affect bubble growth when the temperature and current density are constant. Moreover, they investigated gas bubble (hydrogen and oxygen) dynamics in a single-channel electrolyzer (Li et al., 2019). There, they showed that the bubble detachment diameter is inversely proportional to the flow velocity. Because of the strong influence of electrolyzer geometry on macro scale bubbly flow, Hreiz et al. (2015) studied several different electrochemical configuration sets. In the following work (Abdelouahed et al., 2014), Laser Doppler Velocimetry (LDV) measurements were conducted to observe bubble curtain behavior. Other experimental studies revealed that the bubble curtain velocity is a function of the average electric current (Hine and Murakami, 1980), the mass transfer coefficient is varied for horizontal and vertical electrodes (Fouad and Sedahmed, 1972), and bubble coverage is also a function of current density (Vogt and Balzer, 2005). Vogt (2012) has mentioned that when the bubble coverage tends to unity, the reaction is totally blocked. For the micro-scale, the microphysical phenomena surrounding individual bubbles constitute the overall behavior of the electrolysis process, and consequently, understanding the micro-scale bubble hydrodynamic is vital to reveal the governing mechanisms behind electrolysis. As the studied scale becomes smaller, however, the difficulty of the experiment rises dramatically. Several studies confirm that flow convection in the micro-area is caused by bubble growth, detachment, and coalescence (micro-convection) strongly influences the hydrogen diffusion boundary layer and changes mass transfer (Stephan and Vogt, 1979; Vogt and Balzer, 2005). van der Linde et al. (2018) used a specially designed electrode with pillars and pits to control micro-bubble formation, measuring the bubble radius and attachment force to verify the Fritz radius. Apart from bubble formation and growth, Bashkatov et al. (2024) investigated two modes of coalescence-induced bubble detachment using a dual platinum microelectrode system. By combining high-speed imaging and electrochemical analysis, they demonstrated the importance of bubble-bubble interactions in the departure process. The impact of micro-convection induced by bubble dynamics on mass transfer is significant. Burdyny et al. (2017) demonstrated that micro-convection enhances mass transport primarily by reducing the departure diameter of bubbles from the electrode surface. With advanced visualization techniques, the gas concentration released in the liquid by bubbles can be investigated (Dani et al., 2007; Francois et al., 2011). However, experimental measurements are generally expensive and limited by the available measuring techniques, which usually provide global quantities (e.g., global mass transfer rate, rising velocities), and do not give information about local details, such as small scale multiphase flow field and local mass transfer rate.
@@
Due to the presence of large amounts of gas, conventional optical techniques are unable to detect many critical aspects of multiphase flow fields. Computational Fluid Dynamics (CFD) has become an important tool for the comprehensive study of electrolytic complex multiphase flows (Hawkes et al., 2009; El-Askary et al., 2015). The significant increase in computational power has made direct numerical simulation an important alternative method for studying the detailed dynamics of mass transfer between fixed or deformable interfaces.
@@
The individual bubble growth, multiple bubbles coalescence, and detachment that occur in this region have been numerically studied for a long period. It is confirmed that the diffusion-driven bubble growth dominates the growth stage (Soto, 2019). The bubble detachment radius is approximated by the Fritz radius (Vogt et al., 2004). It can be derived by the force balance acting on the growing and detaching bubbles. Besides, the mass transfer behind the phenomenon still needs to be discussed. van der Linde et al. (2017); van der Linde et al. (2018) simulated the hydrogen concentration field around a growing hydrogen bubble in acidic electrolysis, using a body-fitted axisymmetric finite difference method. Sepahi et al. (2022) employed an Immersed Boundary Method (IBM) to simulate the mass transport for bubbles in gas-evolving electrolysis, indicating that the net transport within the system is governed by the effective buoyancy driving induced by the rising bubble. Very few simulations of deformable-bubble growth, detachment, and rise in the context of electrolysis have been performed using modern sharp-interface methods such as VOF, level set, or front tracking. Moreover, some authors used VOF methods without mass transfer (Lafmejani et al., 2017). Most previous studies, such as those by Sepahi et al. (2022) and Khalighi et al. (2023), assumed undeformable spherical bubbles, neglecting bubble deformation that may arise due to lateral flow during detachment. Simulations of deformable bubbles without the spherical bubble assumption have not been performed very often due to the capillary time-step constraint. Thus, a rigid sphere approximation saves considerable CPU time. Although these models do not apply to the design of real systems, they still provide helpful information and reference solutions for the validation of other models. Recently, Gennari et al. (2022) developed a VOF-based phase-change model for diffusion-driven mass transfer problems using the one-fluid method and a novel algorithm to extrapolate the discontinuous velocity field across the interface to improve interface advection accuracy. This method is applied to study the growth of deformable bubbles on planar electrodes.
In this paper, we investigate all these effects using direct numerical simulation (DNS), which means that we fully resolve all relevant scales of the hydrodynamic and concentration boundary layers. We take advantage of the free code repository Basilisk (http://basilisk.fr/).
@@ 2 Configuration and numerical methods
To simplify the complicated mathematical description of a bubble growing at an electrode in the presence of a surrounding flow, consider a two-phase gas-liquid system represented in fig. 1. We assume a constant room temperature, dilute liquid solutions ($\mathrm{KOH}$, 0.5 mol/L ), and no evaporation of water. Besides, in order to avoid further complications, self-ionisation of water is disregarded due to its low equilibrium constant at room temperature. The equation for the electric potential remains relatively simple. At the cathode, the reaction process is as follows,
@@ 2.1 Problem set-up
The bubble growth and detachment are investigated first under axisymmetry. Due to the geometric axisymmetry, an axisymmetric simulation can also effectively address the problem when studying the growth of a single bubble on a circular electrode in addition to a full 3D simulation. A sketch of the axisymmetric setup and mesh grid is shown in fig. 2. The axisymmetric condition is applied at the left boundary (Z axis), while an outflow boundary condition is set on the top boundary; the other boundaries are treated as no-slip wall. The computational domain is a square. The domain’s characteristic length is $L_{0}=25L_{ref}$, where the reference length is the initial bubble diameter $D_{b}$. The bubble is initialized in simulations with the diameter $D_{b}=0.0127$ mm, The electrode is the flat end of a wire with a diameter of $D_{e}=10D_{b}=0.127$ mm, oriented along the axis of symmetry (Z axis).
As for the more complex configuration, "mutiple nucleation sites", the geomertic axisymmetry does not exist anymore. To address the mutual interaction, a 3D configuration is needed. The multiple nucleation sites are equally spaced at the electrode center, as illustrated in fig. 3.
The initial liquid is set to be saturated, and the saturation ratio is $\zeta=c_{0}/c_{sat}=1$, where $c_{sat}$ is the hydrogen concentration in the saturated liquid and $c_{0}$ is the initial concentration of dissolved gas near the electrode surface. As mentioned in section 1, the production of dissolved gas at the electrode walls creates a locally supersaturated region ($\zeta>1$). It drives the growth of bubbles formed from microscopic pits on the electrode surface due to the heterogeneous nucleation, see Jones et al. (1999); van der Linde et al. (2017). According to the experiment, there should be no bubble present at initial time $t=0$. However, using the VOF method requires the volume fraction of gas need to be initialized. We wait for a nucleation time before computing the volume change. The bubble size is fixed during this stage ($t<t_{n}$), while the concentration of dissolved hydrogen continues to increase. This approach enables the development of a concentration field around the bubble by the time nucleation occurs, better reflecting experimental observation. The nucleation time varies due to many factors, like the electrode material and the current density. In present work, we set the nucleation time $t_{n}=0.02$ s, which is observed from Glas and Westwater (1964). The control parameters for the electrolytically generated bubbly flow are the cathodic current density, and the contact angles for different wettability of the electrode surface. The current density $I$ can give the molar flux of hydrogen ($\mathrm{H_{2}}$) by Faraday’s law,
@@
where $I=i/A$, $i$ is the total electric current, $A$ is the cross-section area ($A=\pi/4D_{e}^{2}$); $F$ is Faraday’s constant ($F=96485.3$ As/mol). To account for the flux of H2 across the active area of the electrode, a Neumann boundary condition for the gas concentration is applied to the electrode wall (r-axis for axis-symmetric simulation, XY plane for 3D simulation),
@@
The boundary influx from the bubble-free region directly leads to an increase of the local concentration of dissolved hydrogen, leading to the development of a hydrogen concentration layer. This process facilitates hydrogen bubble formation, driving its growth and detachment. Simulations performed in the present work are classified into two groups. The configurations are listed in Table 1. The axisymmetric configuration is mainly used for numerical method verification by comparing the bubble growth with analytical and experimental results. The 3D configuration is designed to study the bubble coalescence and detachment.
@@
Single bubble (Axisymmetric) $I(Am^{-2})$
No. $\theta\degree$ $z\times r(L_{ref})$ mesh level 100
1 90 $25\times 25$ 8 300
2 90 $25\times 25$ 9 1000
3 90 $25\times 25$ 10 -
4 35 $25\times 25$ 8 -
5 35 $25\times 25$ 9 -
6 35 $25\times 25$ 10 -
3D
No. $\theta\degree$ $x\times y\times z(L_{ref})$ nucleation sites -
7 90 $25\times 25\times 25$ 1 -
8 90 $25\times 25\times 25$ 2 -
9 90 $25\times 25\times 25$ 4 -
10 35 $25\times 25\times 25$ 1 -
11 35 $25\times 25\times 25$ 2 -
12 35 $25\times 25\times 25$ 4 -
Table 1: Simulation parameters for cases with a single bubble and mutiple bubble nucleation sites. All the simulation are performed at 3 different current densities, as listed in the last column.
The physical parameters remain the same for all the simulation cases (see Table 2). The only exception is that gravity acceleration is set to zero for the axisymmetric configuration. To reduce the computational cost due to the time step limitation of the surface tension scheme, $\sigma$ is decreased by a factor of $10^{-4}$. Besides, the density of hydrogen is relatively small, and the large density ratio for this system ($\rho_{c}/\rho_{d}=12450$) would be a problem for the numerical simulation. It is observed that a large density ratio will slow down the convergence of the multigrid solver. To circumvent this problem, the hydrogen density is increased to $\rho_{d}=0.8\,\text{kg/m}^{3}$, the molar mass in increased by the same factor. so the volume change is not influenced ($\Delta V\propto M/\rho_{d}$). The zero gravity settings are only applied for the axisymmetric configuration. When it comes to the investigation of bubble detachment, gravity acceleration is applied at the direction of - Z in 3D configuration with the value of $9.8\mathrm{m/s^{2}}$.
@@
$\rho_{c}$ Electrolyte density 996 [$\mathrm{kg/m^{3}}$]
$\rho_{d}$ Hydrogen density 0.8 [$\mathrm{kg/m^{3}}$]
$g$ Gravity acceleration 9.8 [$\mathrm{m/s^{2}}$]
$M$ Molar mass of hydrogen $2.0\times 10^{-2}$ [$\mathrm{kg/mol}$]
$c_{0}$ H2 initial concentration 0.002 [$\mathrm{mol/m^{3}}$]
$c_{\text{sat}}$ H2 saturated concentration 0.02 [$\mathrm{mol/m^{3}}$]
$\mu_{c}$ Electrolyte viscosity $8.32\times 10^{-4}$ [$\mathrm{kg/(m\cdot s)}$]
$\mu_{d}$ Hydrogen viscosity $8.96\times 10^{-6}$ [$\mathrm{kg/(m\cdot s)}$]
$\nu_{c}$ Electrolyte kinematic viscosity $8.35\times 10^{-7}$ [$\mathrm{m^{2}/s}$]
$\sigma$ Surface tension $7.5\times 10^{-2}$ [$\mathrm{N/m}$]
$F$ Faraday’s constant $96485.3$ [$\mathrm{C/mol}$]
$D$ Hydrogen diffusion coefficient $7.38\times 10^{-9}$ [$\mathrm{m^{2}/s}$]
$L_{\text{ref}}$ Reference length $1.27\times 10^{-5}$ [$\mathrm{m}$]
$D_{b}$ Initial bubble diameter $1.27\times 10^{-5}$ [$\mathrm{m}$]
$D_{e}$ Electrode diameter $1.27\times 10^{-4}$ [$\mathrm{m}$]

### v4

@@ Abstract
Three-dimensional direct numerical simulation of electrolysis is applied to investigate the growth and detachment of bubbles at electrodes. The moving gas-liquid interface is modeled employing the VOF-based method. To ensure the accuracy of the simulations, a mesh-independence study has been performed. The simulations include the growth phase of the bubbles, followed by their detachment from the electrode surface, and the results are validated with analytical models and experimental data. The bubble growth is diffusion-controlled, leading to the scaling $R\propto t^{1/2}$, but our simulation overpredicts the growth exponent during the initial stage. We further demonstrate that the number of nucleation sites significantly affects gas transport, as quantified by the Sherwood number. The influences of contact angle and nucleation site on bubble detachment are also examined. The predicted detachment radius varies linearly with contact angle, consistent with Fritz’s linear relation between the volume-equivalent radius and contact angle, confirming that the surface tension is the dominant attachment force. Finally, as the nucleation sites increase, the induced bubble coalescence accelerates the bubble detachment. Taken together, these findings give us valuable insights into improving gas bubble removal and enhancing overall electrolysis efficiency.
@@ 1 Introduction
The production of green hydrogen through water electrolysis is expected to be an important technology in achieving global net-zero emissions (Turner2004; Holladay2009; Dawood2020). However, it is slow and inefficient in many situations. The attached bubbles in such electrochemical devices reduce the efficiency of electrolyzer systems by blocking the active electrode sites or by increasing the ohmic resistance (Swiegers2021). Maintaining a bubble-free electrode surface is therefore crucial for highly efficient $\mathrm{H_{2}}$ production. Consequently, a detailed understanding of bubble evolution dynamics is highly desirable for developing strategies to enhance water electrolysis efficiency. Analytical solutions for this problem are not generally available, and only a few exact solutions can be derived using extremely simplified assumptions (Epstein1950; Scriven1959). However, they can still serve as valuable references for experimental and numerical studies (Glas1964; Brandon1985; dapkus1986; vanderlinde2017a; Taqieddin2017; Soto2018; Angulo2020; Zhang2023).
Several experiments have confirmed that bubble growth is primarily governed by mass transport arising from gradients in the dissolved hydrogen concentration. Glas1964 performed a fundamental experiment for hydrogen bubbles generated on a flat electrode. The experimental findings demonstrate that the asymptotic growth of electrochemically generated bubbles follows the same functional relationship given by the analytical solution for a suspended bubble growing in a supersaturated liquid, see Scriven1959. Recent studies have already investigated this problem on different scales. Macroscopic processes are mainly convection phenomena in the electrolyte solution, which is observed at the electrolytic cell scale. li2018 visualized experimentally on two-phase flow at the anode side of a proton exchange membrane electrolyzer. They concluded that the inlet velocity does not affect bubble growth when the temperature and current density are constant. Furthermore, they examined hydrogen and oxygen bubble dynamics in a single-channel electrolyzer (li2019), revealing that the bubble detachment diameter varies inversely with flow velocity. Given the significant influence of electrolyzer geometry on macroscale bubbly flow, hreiz2015 analyzed a variety of electrochemical configurations.
@@
Abdelouahed2014 conducted Laser Doppler Velocimetry (LDV) measurements to observe the behavior of the bubble curtain. Other experimental studies revealed that the bubble curtain velocity is a function of the average electric current (hine1980), the mass transfer coefficient is varied for horizontal and vertical electrodes (fouad1972), and bubble coverage is also a function of current density (vogt2005). vogt2012 has mentioned that when the bubble coverage approaches unity, the reaction is totally blocked. For the micro-scale, the micro-physical phenomena surrounding individual bubbles govern the overall behavior of the electrolysis process, and consequently, understanding the micro-scale bubble hydrodynamics is vital to reveal the governing mechanisms behind electrolysis. As the studied scale becomes smaller, however, the difficulty of the experimentation increases dramatically. Several studies confirm that flow convection in the micro-area is caused by bubble growth, detachment, and coalescence. Such micro-convection strongly influences the hydrogen diffusion boundary layer and changes mass transfer (Stephan1979; vogt2005). For example, Linde2018 used a specially designed electrode with pillars and pits to control micro-bubble formation, measuring the bubble radius and attachment force to verify the Fritz radius. Apart from bubble formation and growth, Bashkatov2024 investigated two modes of coalescence-induced bubble detachment using a dual platinum microelectrode system. By combining high-speed imaging and electrochemical analysis, they demonstrated the importance of bubble-bubble interactions in the departure process. The impact of micro-convection induced by bubble dynamics on mass transfer is significant. Burdyny2017 demonstrated that micro-convection enhances mass transport primarily by reducing the departure diameter of bubbles from the electrode surface. With advanced visualization techniques, the gas concentration released in the liquid by bubbles can be investigated (dani2007; francois2011). However, experimental measurements are generally expensive and constrained by the available measuring techniques, which usually provide global quantities (e.g., global mass transfer rate, rising velocities), and do not give information about local details, such as small-scale multiphase flow field and local mass transfer rate.
@@
Due to the presence of large amounts of gas, conventional optical techniques are unable to detect many critical aspects of multiphase flow fields. Computational Fluid Dynamics (CFD) has become an important tool for the comprehensive study of electrolytic complex multiphase flows (Hawkes2009; El-Askary2015). The significant increase in computational power has made direct numerical simulation an important alternative method for studying the detailed dynamics of mass transfer between fixed or deformable interfaces.
@@
The individual bubble growth, multiple bubbles coalescence, and detachment that occur in this region have been numerically studied for a long period. It is confirmed that the diffusion-driven bubble growth dominates the growth stage (Soto2018). The bubble detachment radius is approximated by the Fritz radius (vogt2004). It can be derived from the force balance acting on the growing and detaching bubbles. Besides, the mass transfer mechanisms underlying this phenomenon require further discussion. vanderlinde2017a; Linde2018 simulated the hydrogen concentration field around a growing hydrogen bubble in acidic electrolysis, using a body-fitted axisymmetric finite difference method. sepahi2022a employed an Immersed Boundary Method (IBM) to simulate the mass transport for bubbles in gas-evolving electrolysis, indicating that the net transport within the system is governed by the effective buoyancy driving induced by the rising bubble. Very few simulations of deformable-bubble growth, detachment, and rise in the context of electrolysis have been performed using modern sharp-interface methods such as VOF, level set, or front tracking. Moreover, some authors used VOF methods without mass transfer (lafmejani2017). Most previous studies, such as those by sepahi2022a and khalighi2023, assumed undeformable spherical bubbles, neglecting bubble deformation that may arise due to lateral flow during detachment. Due to the time step limitations imposed by capillary effects, employing a rigid sphere approximation considerably reduces CPU time. Although these models are not suitable for designing actual systems, they still provide useful information and reference solutions for validating novel theories. Thus, simulations of bubble growth in electrolysis without the spherical bubble assumption have rarely been performed. Recently, Gennari2022 proposed a VOF-based phase-change model for diffusion-driven mass transfer problems using the one-fluid method and a novel algorithm to extrapolate the discontinuous velocity field across the interface to improve interface advection accuracy. This method is applied to study the growth of deformable bubbles on planar electrodes.
This study employs direct numerical simulation (DNS) to investigate all the aforementioned effects, meaning that we have fully resolved all relevant scales of the hydrodynamic and concentration boundary layers. Compared to existing numerical studies, we are the first to utilize three-dimensional DNS to simulate the deformable interface motion of hydrogen bubbles at the electrode. The fully resolved flow field provides crucial insight into the mechanisms of bubble growth, interaction, and detachment. All simulations are performed using the free code repository Basilisk (http://basilisk.fr/).
@@ 2 Configuration and numerical methods
To simplify the complicated mathematical description of a bubble growing at an electrode in the presence of a surrounding flow, consider a two-phase gas-liquid system represented in Fig.˜1. We assume a constant room temperature, dilute liquid solutions ($\mathrm{KOH}$, $0.5\mathrm{mol/L}$), and no evaporation of water. Besides, in order to avoid further complications, self-ionization of water is disregarded due to its low equilibrium constant at room temperature. The equation for the electric potential remains relatively simple. At the cathode, the reaction process is as follows,
@@ 2.1 Problem set-up
We start with axisymmetric simulations to validate the numerical mass transfer method provided by Gennari2022. Due to the geometric axisymmetric nature of the single bubble attached to a circular electrode, axisymmetric modeling not only effectively resolves the problem but also reduces computational costs while achieving a three-dimensional simulation. A sketch of the axisymmetric setup and mesh grid is shown in Fig.˜2. The symmetry axis is along $z$, while an outflow boundary condition is set on the top boundary. The other boundaries are treated as no-slip walls. The computational domain is a square. The domain’s size is $L_{0}=25D_{b}$, where $D_{b}$ is the initial bubble diameter. The bubble is initialized in simulations with the diameter $D_{b}=0.0127$ mm, The electrode is the flat end of a wire with a diameter of $D_{e}=10D_{b}=0.127$ mm, oriented along the radial axis (r axis).
As for the more complex case of multiple nucleation sites, axial symmetry no longer holds. To address the mutual interaction between bubbles, a 3D configuration is required. The multiple nucleation sites are equally spaced at the electrode center, as illustrated in Fig.˜3.
The initial liquid is set to be saturated, and the saturation ratio is $\zeta=c_{0}/c_{s}=1$, where $c_{s}$ is the hydrogen concentration in the saturated liquid and $c_{0}$ is the initial concentration of dissolved gas near the electrode surface. As mentioned in Section˜1, the production of dissolved gas at the electrode walls creates a locally supersaturated region ($\zeta>1$). It drives the growth of bubbles formed from microscopic pits on the electrode surface due to the heterogeneous nucleation, see Jones1999; vanderlinde2017a. According to the experiment, there should be no bubble present at the initial time $t=0$. However, using the VOF method requires the volume fraction of gas to be initialized. We wait for a nucleation time before computing the volume change. The bubble size is fixed during this stage ($t<t_{n}$), while the concentration of dissolved hydrogen continues to increase. This approach enables the development of a concentration field around the bubble by the time nucleation occurs, better reflecting experimental observation. The nucleation time varies due to many factors, like the electrode material and the current density. In the present work, we set the nucleation time $t_{n}=0.02\,\mathrm{s}$, a value derived from experimental work(Glas1964). The control parameters for the electrolytically generated bubbly flow are the cathodic current density and the contact angles for different wettability of the electrode surface. The current density $I$ can give the molar flux of hydrogen ($\mathrm{H_{2}}$) by Faraday’s law,
@@
where $I=i/A$, $i$ is the total electric current, $A$ is the cross-section area ($A=\pi/4D_{e}^{2}$); $F$ is Faraday’s constant ($F=96485.3$ As/mol). To account for the flux of H2 across the active area of the electrode, a Neumann boundary condition for the gas concentration is applied to the electrode wall (r-axis for axis-symmetric simulation, xy plane for three-dimensional simulation),
@@
The boundary influx from the bubble-free region directly leads to an increase in the local concentration of dissolved hydrogen, leading to the development of a hydrogen concentration layer. This process facilitates hydrogen bubble formation, driving its growth and detachment. Simulations performed in the present work are classified into two groups. The configurations are listed in Table˜1. The axisymmetric configuration is primarily employed for numerical verification and sensitivity analysis, whereas the 3D configuration is used for most of the simulations on bubble growth evolution.
@@
Single bubble (Axisymmetric)
No. $\theta\degree$ $z\times r(D_{b})$ mesh level
1 35–90 (sweep: 35, 40, …, 90) $25\times 25$ 8
2 90(hydrophobic) $25\times 25$ 9
3 90(hydrophobic) $25\times 25$ 10
4 35(hydrophilic) $25\times 25$ 9
5 35(hydrophilic) $25\times 25$ 10
Three-dimensional
No. $\theta\degree$ $x\times y\times z(D_{b})$ nucleation sites
6 35–90 (sweep: 35, 40, …, 90) $25\times 25\times 25$ 1
7 90(hydrophobic) $25\times 25\times 25$ 2
8 90(hydrophobic) $25\times 25\times 25$ 4
9 35(hydrophilic) $25\times 25\times 25$ 2
10 35(hydrophilic) $25\times 25\times 25$ 4
Table 1: Configurations for the three-dimensional and axisymmetric simulations.
The physical parameters remain the same for all the simulation cases (see Table˜2). Gravity acceleration is applied in the z-z direction with the value of $9.8\mathrm{m/s^{2}}$.
@@
$\rho_{c}$ Electrolyte density 996 $\mathrm{kg/m^{3}}$]
$\rho_{d}$ Hydrogen density 0.8 $\mathrm{kg/m^{3}}$]
$g$ Gravity acceleration 9.8 $\mathrm{m/s^{2}}$]
$M$ Molar mass of hydrogen 0.02 $\mathrm{kg/mol}$]
$c_{0}$ H2 initial concentration 0.002 $\mathrm{mol/m^{3}}$]
$c_{s}$ H2 saturated concentration 0.02 $\mathrm{mol/m^{3}}$]
$\mu_{c}$ Electrolyte viscosity $8.32\times 10^{-4}$ $\mathrm{kg/(m\cdot s)}$]
$\mu_{d}$ Hydrogen viscosity $8.96\times 10^{-6}$ $\mathrm{kg/(m\cdot s)}$]
$\nu_{c}$ Electrolyte kinematic viscosity $8.35\times 10^{-7}$ $\mathrm{m^{2}/s}$]
$\sigma$ Surface tension $7.5\times 10^{-6}$ $\mathrm{N/m}$]
$F$ Faraday’s constant $96485.3$ $\mathrm{C/mol}$]
$D$ Hydrogen diffusion coefficient $7.38\times 10^{-9}$ $\mathrm{m^{2}/s}$]
$D_{b}$ Initial bubble diameter $1.27\times 10^{-5}$ $\mathrm{m}$]
$R_{0}$ Initial bubble radius $6.35\times 10^{-6}$ $\mathrm{m}$]
$D_{e}$ Electrode diameter $1.27\times 10^{-4}$ $\mathrm{m}$]
$t_{n}$ Bubble nucleation time $0.02$ $\mathrm{s}$]
$P_{0}$ Ambient pressure $1.01325\times 10^{5}$ $\mathrm{N/m^{2}}$]
$T_{0}$ Ambient temperature $298.15$ $\mathrm{K}$]
$\mathcal{R}$ Universal gas constant $8.3145$ $\mathrm{J\cdotp mol^{-1}\cdotp K^{-1}}$]
$H_{s}^{cp}$ Henry’s law solubility constant for $\mathrm{H_{2}}$ $7.5\times 10^{-6}$ $\mathrm{mol/(m^{3}Pa)}$]
$P$ Henry’s law partition coefficient 53.3 -